OpenStudy (firejay5):

Show work!!!! :DDD Medal will be rewarded!! :D 6. (1 over 3x - 2) + (5 over x) = 0

OpenStudy (firejay5):

6. \[\frac{ 1 }{ 3x - 2 } + \frac{ 5 }{ x } = 0\]

OpenStudy (anonymous):

First you need to make the fractions have the same denominator: \[\frac{ x }{ x(3x -2) } + \frac{ 5(3x - 2) }{ x(3x-2) } = 0\] Then just add the fractions together (x + 5(3x - 2) becomes the numerator) and simplify the numerator. From there, you know that the numerator will equal zero if the entire expression equals zero.

OpenStudy (anonymous):

^What he saidm and here are a few values x cannot have, x=0, and x=2/3. This is so because at these values, the function will be non existent. (case being of denominator=0)

OpenStudy (firejay5):

so @Fruitbasket will it be 5x(3x - 2) = 0

OpenStudy (anonymous):

^No, what he said, si the quation is x+15x-10=0 That is 16x=10 i.e. x=5/8

OpenStudy (anonymous):

To which I added the two values x cannot have and the reason for me stating so.

OpenStudy (firejay5):

I need help not a lecture, did I do it right

OpenStudy (anonymous):

No. The "lecture" was just a solution in detail to make it clear.

OpenStudy (anonymous):

I think its x ------- 15x - 10

OpenStudy (anonymous):

which goes 15x - 10x but im not 100%

OpenStudy (firejay5):

How did you get that

OpenStudy (firejay5):

15x^2 - 10

OpenStudy (anonymous):

15x-10 =5x 5x=0 - - 5 5 x=0

OpenStudy (anonymous):

I think its x ------- =0 multiply 1x to everything 15x - 10 15x - 10x =0 now combine like terms 5x=0 now get the variable by its self by dividing 5 - - 5 5 x=0 which gives you X = 0

OpenStudy (anonymous):

if I helped and its correct give a best answer for people to know hehe was I any help

OpenStudy (anonymous):

do you understand what I put

OpenStudy (firejay5):

@michaelc96 should it be 15x^2 - 10x

OpenStudy (firejay5):

5x(3x - 2) = 0 divide by 5 x( 3x - 2) = 0 Split into 2 problems x = 0 or 3x - 2 = 0 <--- x = 2/3

OpenStudy (firejay5):

right?

OpenStudy (anonymous):

so your saying should 15x be to the second power

OpenStudy (firejay5):

yes

OpenStudy (anonymous):

I would say yes imam see if @Mertsj can help

OpenStudy (firejay5):

the answer should be x = 5/8

OpenStudy (mertsj):

Right away, before we do anything, we see that x cannot be 2/3 or 0 Now we can go ahead and write both fractions with the common denominator x(3x-2)

OpenStudy (mertsj):

Well actually, I didn't see that it was an equation so we should just multiply every term on both sides by the common denominator. This will eliminate all denominators and we will end up with:

OpenStudy (mertsj):

\[x(3x-2)\times\frac{1}{3x-2}+x(3x-2)\times \frac{5}{x}=x(3x-2)\times 0\]

OpenStudy (mertsj):

Which, cancelling common factors gives us: \[x+5(3x-2)=0\]

OpenStudy (mertsj):

\[x+15x-10=0\]

OpenStudy (mertsj):

\[16x-10=0\]

OpenStudy (mertsj):

\[16x=10\]

OpenStudy (mertsj):

\[x=\frac{10}{16}=\frac{5}{8}\]

OpenStudy (anonymous):

man you know your stuff hehe

OpenStudy (firejay5):

that's why I refer to him! :D

OpenStudy (mertsj):

ty

OpenStudy (anonymous):

pfft yet I was the on to call him im guessing your always in this part

OpenStudy (mertsj):

Whatever, dude

OpenStudy (mertsj):

Just so the problem got solved. Everyone's happy, right?

OpenStudy (anonymous):

yup

OpenStudy (firejay5):

yep